OpenStudy (anonymous):

Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form. my answer is y=-x-2

OpenStudy (wasiqss):

gradient is -1

OpenStudy (wasiqss):

plug in points and get the equation y=mx+c 2=4+c c=-2 hence y=-x+2

OpenStudy (wasiqss):

i mean y=-x-2

OpenStudy (anonymous):

thank u wasiqss

OpenStudy (wasiqss):

np :)